eviant:

The domain for f(x) and g(x) is all real numbers. Let f(x) = 3x^2 + x-7 and g(x) = x^2 + 4x-5 Find (f-g)(x).

eviant:

@dude

Nnesha:

\[\rm (f-g)(x)=f(x)-g(x)\]

Nnesha:

\[\rm \color{Red}{f(x) = 3x^2 + x-7}, ~~~\color{blue}{g(x) = x^2 + 4x-5}\] \[\rm (f-g)(x)=\color{Red}{f(x)}-\color{blue}{g(x)}\] \[\rm (f-g)(x)=(\color{Red}{3x^2 + x-7})-(\color{blue}{ x^2 + 4x-5})\] 1) distribute the negative sign 2) combine like terms

Nnesha:

\[\rm (f-g)(x)=(\color{Red}{3x^2 + x-7})-(\color{blue}{ x^2 + 4x-5})\] is same as \[\rm (f-g)(x)=(\color{Red}{3x^2 + x-7})-\color{orange}{1}(\color{blue}{ x^2 + 4x-5})\]

eviant:

@Nnesha, thanks I got it, I think

eviant:

I got 3x^2-3x-2

Nnesha:

how?

eviant:

I subtracted 3x^2-x^2, which game 2x^2

eviant:

then, I subtracted x-4x which equals -3x

Nnesha:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @eviant I subtracted 3x^2-x^2, which game 2x^2 \(\color{#0cbb34}{\text{End of Quote}}\) yes correct so it should be 2x^2

eviant:

finally, I did -7+5, which gave me -2

Nnesha:

good job!

eviant:

@Nnesha is it okay to add the last part?

eviant:

the -7+5 part?

eviant:

@dude @AngeI

Nnesha:

it will NOT be okay if you don't add.

Nnesha:

that's part of the answer "the constant" term

Nnesha:

(f-g)(x)= `2x^2-3x-2`

eviant:

@Nnesha so its correct, right?

Nnesha:

yes